A piulley supporting a load of 400 N is mounted at C on a horizontal beam .if the beam weighs 800N,find the support reactions at A and B . neglect the weights of the pulley.

1
Expert's answer
2022-01-27T22:56:02-0500

Assume pully at center C

Total load on Beam at "W_c = 400 + 800 = 1200 N"


Support reactions at "R_A = \frac{W_c}{2}= \frac{1200}{2}=600 N"


Support reactions at "R_B = W_C-R_A = 1200-600 = 600 N"